OpenStudy (anonymous):

Find a formula for a quadratic function with vertex at (-1, 2) that passes through the point (0,4)

OpenStudy (anonymous):

vertex form is y = a(x-h)^2 + k so I have y = a(x-(-1))^2 - 2 4 = a(0-(-1))^2 - 2 is this correct?